The time it takes a randomly selected employee to perform a task is normally distributed with a mean value of 120 seconds and a standard deviation of 20 seconds.

Calculate the probability that a randomly selected employee will complete:

1. The task within 100 to 130 seconds

2. In more than 75 seconds


Question 2

The table below shows whether students are left handed or right handed by sex. The total number of students who were interviewed was 100. The responses was as follows;


Male Female Total

Righ habded 38 42 80

Left handed 12 8 20

Total 50 50 100


1. What is the probability of selecting a left-handed male?

2. What is the probability of selecting a right-handed female?


1
Expert's answer
2021-05-02T09:48:46-0400

Question 1

"\mu = 120 \\

\sigma = 20"

1.

"P(100<X<130) = P(\frac{100-120}{20}<Z<\frac{130-120}{20}) \\

= P(-1<Z<0.5) \\

= P(Z<0.5) – P(Z< -1) \\

= 0.6914 -0.1586 \\

= 0.5328"

2.

"P(X>75) = 1 -P(X<75) \\

= 1 -P(Z<\frac{75-120}{20}) \\

= 1 -P(Z< -2.25) \\

= 1 -0.0122 \\

= 0.9878"

Question 2



1. P(left-handed male) "= \frac{Left-hunded \;male}{Total}"

"= \frac{8}{100} \\

= 0.08"

2. P(right-handed female) "= \frac{Right-hunded \;female}{Total}"

"= \frac{42}{100} \\

= 0.42"